Data sets A and B are dependent. Find d. A4.5 5.5 7.4 4.4 4.5 B| 6.9 5.8 5.7 5.6 7 OA. 0.89 Tom B. $4.94 C. -0.76 OD. 0.58
Answer #1

First we will find d=A-B

A B d
4.5 6.9 -2.4
5.5 5.8 -0.3
7.4 5.7 1.7
4.4 5.6 -1.2
4.5 7 -2.5
Total -4.7

So \overline{d}=\frac{\sum d}{n}=\frac{-4.7}{5}=-0.94

Hence answer here is B. -0.94